### v2

# Cortex AISQL: A Production SQL Engine for Unstructured Data

Abstract.
Integrating AI’s semantic capabilities directly into SQL is a key goal, as it would enable users to write declarative queries that seamlessly blend relational operations with semantic reasoning. This would allow for effortless querying of both structured and unstructured data. However, achieving production-scale efficiency for semantic operations presents significant hurdles. Semantic operations inherently cost more than traditional SQL operations and possess different latency and throughput characteristics, making their application at large scales prohibitively expensive. Moreover, current query execution engines are not designed to optimize semantic operations, often leading to suboptimal query plans. This optimization challenge is compounded because the cost and selectivity of semantic operations are typically unknown at the time of compilation. Snowflake’s Cortex AISQL is a query execution engine that addresses these challenges through three novel techniques informed by production deployment data from Snowflake customers. First, AI-aware query optimization treats AI inference cost as a first-class optimization objective, reasoning about large language model (LLM) cost directly during query planning to achieve 2–8 speedups. Second, adaptive model cascades reduce inference costs by routing most rows through a fast proxy model while escalating uncertain cases to a powerful oracle model, achieving 2–6 speedups while maintaining 90–95% of oracle model quality. Third, semantic join query rewriting lowers the quadratic time complexity of join operations to linear through reformulation as multi-label classification tasks, achieving 15–70 speedups with often improved prediction quality. AISQL is deployed in production at Snowflake, where it powers diverse customer workloads across analytics, search, and content understanding.
1. Introduction
The growth of large language models (LLMs) (brown2020language; ouyang2022training; touvron2023llama; wei2022chain) has transformed how organizations interact with data. Modern enterprises now store vast volumes of unstructured content – documents, images, audio, and text – alongside traditional structured data in relational tables. While contemporary data warehouses excel at processing structured data with SQL, they lack native support for semantic reasoning over unstructured information. As a result, users who need to filter documents by topic relevance, join tables based on semantic similarity, or summarize customer feedback must export data to external systems, write custom scripts, or build complex pipelines that orchestrate LLM APIs outside the database. The result is significant friction: data must be moved across systems, increasing latency and cost; manual orchestration becomes error-prone; and the database loses the ability to optimize the workload end-to-end.
We present Snowflake’s Cortex AISQL, a SQL engine that enables users to talk to their data, both structured and unstructured, through native AI operators integrated directly into SQL. AISQL extends SQL with primitive operators that bring LLM capabilities to the relational model: AI_COMPLETE for text generation, AI_FILTER for semantic filtering, AI_JOIN for semantic joins, AI_CLASSIFY for categorization, and AI_AGG and AI_SUMMARIZE_AGG for semantic aggregations and text summarizations, respectively. These operators naturally compose with traditional SQL constructs, allowing users to write declarative queries that blend relational operations with semantic reasoning. For example, a product manager analyzing customer feedback can filter support transcripts for conversations where customers expressed frustration (AI_FILTER), semantically join them with a product catalog to identify which specific products were discussed (AI_JOIN), classify the severity of each issue (AI_CLASSIFY), and generate executive summaries grouped by product category (AI_SUMMARIZE_AGG) – all within a single declarative SQL query that combines structured sales data with unstructured conversational text.
AISQL builds on an emerging body of work in semantic query processing. Systems such as LOTUS (patel2025semantic) introduced semantic operators for processing pandas-like dataframes, Palimpzest (liu2025palimpzest) provided a declarative framework for LLM-powered data transformations, while ThalamusDB (jo2025thalamusdb) explored multimodal query processing with natural language predicates. Commercial systems have also adopted this direction: Google BigQuery (fernandes2015bigquery) provides AI functions such as AI.CLASSIFY for semantic operations, while Azure SQL exposes row-level LLM invocations within SQL queries. However, making semantic operators efficient at production scale in a distributed database remains a fundamental challenge. Such operators are orders of magnitude more expensive than traditional SQL operations. For example, a single AI_FILTER predicate applied to a million-row table could invoke an LLM thousands of times, leading to prohibitive execution costs and latency. Traditional query optimization heuristics, such as minimizing join costs or pushing filters below joins, can produce extremely inefficient plans when AI operators are involved. Furthermore, semantic operators behave as black boxes to the query optimizer. Unlike conventional predicates, their cost and selectivity cannot be inferred from historical statistics or column distributions. Estimating how many rows will pass an AI-based filter, or how expensive that evaluation will be, is nearly impossible at compile time. Yet, these placement and cost estimates can alter overall query performance by orders of magnitude.
The AISQL query execution engine addresses these challenges through a combination of novel optimization techniques and runtime adaptations, showcasing that semantic operations can be effectively integrated into database engines, thus enabling use cases that analyze both structured and unstructured data. Production deployment data from Snowflake customers directly shaped our design choices. We observed that LLM inference costs far exceed traditional query processing expenses, with AI operations accounting for the dominant share of execution budgets. Multi-table queries are also prevalent, representing nearly 40% of all AISQL workloads and consuming well over half of aggregate computation time (see Section 4 for details). These observations guide the design of our optimization strategies that are applied at both the query and individual operator level:
AI-aware query optimization. The AISQL optimizer treats AI inference cost as a first-class optimization objective. Rather than optimizing solely for traditional metrics like join cardinality, the optimizer considers the monetary and computational cost of LLM invocations when determining operator placement and predicate evaluation order. For instance, the optimizer may pull expensive multi-modal AI_FILTER predicates above joins or reorder multiple AI predicates based on their relative cost, even when traditional selectivity-based heuristics would suggest otherwise. These optimizations enhance the performance of the AISQL query, with the degree of improvement depending on factors such as the selectivities of various query operators (e.g., predicates, joins). Across synthetic datasets and queries modeled after actual Snowflake customer workloads, we have measured 2–8 more performant execution plans. During execution, runtime statistics about predicate cost and selectivity enable dynamic reordering to adapt to actual data distributions. By reasoning about LLM cost directly during query planning, AISQL now makes even complex multi-modal workloads efficient and scalable.
Adaptive model cascades. To reduce inference costs and maintain query prediction accuracy, AISQL implements model cascades for AI_FILTER operations. A fast and inexpensive proxy model processes most rows, while a powerful oracle model handles only uncertain cases. The system learns confidence thresholds online through importance sampling and adaptive threshold estimation, partitioning rows into accept, reject, and uncertainty regions. Our algorithm is tailored for a distributed execution environment and achieves 2–6 speedups across various benchmarks while maintaining 90–95% of oracle model quality. On the NQ dataset, the cascade achieves 5.85 speedup and preserves the near-identical F1 score; see Section 6.2 for details.
Query rewriting for semantic joins. Naive semantic joins using AI_FILTER typically require a quadratic number of inference calls making them impractical for large datasets. The AISQL engine rewrites certain join patterns into multi-label classification problems using AI_CLASSIFY, reducing complexity from quadratic to linear. Our implementation not only achieves 15–70 speedups across a number of datasets but often improves prediction quality by enabling better comparative reasoning. On the CNN dataset, the rewrite achieves 69.52 speedup, reducing execution time from 4.4 hours to 3.8 minutes; see Section 6.3 for details.
AISQL is deployed in production at Snowflake, where it powers diverse customer workloads across analytics, search, and content understanding. It is fully integrated into the SQL engine and is optimized for efficiency and high prediction quality, as demonstrated by our extensive experimental analysis.
The remainder of this paper is organized as follows. Section 2 describes Snowflake’s architecture with the Cortex Platform for AI workloads. Section 3 introduces AISQL’s operator semantics and SQL interface. Section 4 analyzes customer workload characteristics. Section 5 presents the query execution engine, including AI-aware query optimization, adaptive model cascades, and query rewriting for semantic joins. Section 6 provides an experimental evaluation. Section 7 discusses related work. We conclude in Section 8.
2. Architecture
In this section, we describe how Snowflake’s architecture has evolved to support AI workloads. Separation of compute and storage is a key design decision in Snowflake’s architecture that lets users scale query-processing resources (i.e., virtual warehouses) independently of the amount of data stored in Snowflake tables. The Cloud Services layer handles authentication, query compilation, and coordination of SQL execution.
| Operator | Function Signature | Description | Multimodal |
| Project each row to text based on task instruction | ✓ | ||
| Project each row to boolean based on predicate | ✓ | ||
| Select indices from that satisfy | ✓ | ||
| Project each row to discrete category from candidate set per instruction | |||
| Reduce a column to text aggregate based on the instruction | |||
| AI_SUMMARIZE_AGG | Reduce a column to a text summary |
The commoditization of AI, driven by the rapid growth and adoption of LLMs, has introduced new workloads and reshaped existing data-management tasks. AI agents today need to operate on data traditionally stored and managed in a warehouse. At the same time, users want to run analytical workloads and extract insights from unstructured data (e.g., documents, images, audio, and video). Section 4 analyzes AI workloads at Snowflake.
To support these workloads, we introduced the Cortex Platform into the Snowflake architecture (Figure 1). The Cortex Platform is a multi-tenant service that executes both interactive (e.g., via REST) and batch AI workloads (e.g., via SQL). Its primary components are (a) Inference Engines, (b) the Scheduler, and (c) the API Service.
The Inference Engines provide the core capabilities of the Cortex Platform. Each engine is a specialized service that runs trained models (e.g., LLMs) to generate predictions or insights from data. An engine manages both the underlying hardware (e.g., GPUs) and the inference stack (e.g., vLLM). The Cortex Platform automatically scales engines up or down to match fluctuations in inference demand.
The Scheduler is the component responsible for orchestrating requests and assigning them to the most appropriate Inference Engine. For instance, an inference request for a specific LLM (e.g., Llama 3.1 70B) is routed to an engine that already hosts that model and is ready to serve it.
The API Service acts as the front-end for the Cortex Platform. It receives inference requests from either the Cloud Services layer or the Query Processing layer, applies API-specific business logic, and forwards the request to the Scheduler.
The Cortex Platform supports a broad set of models from partners including OpenAI, Anthropic, and Meta, enabling users to build and evaluate AI applications easily. For each inference request, the Cortex Platform determines whether to execute it using one of its Inference Engines or forward it to a partner endpoint. For instance, requests for GPT models are routed to OpenAI’s endpoints.
3. AISQL Operators
In this section, we introduce the core operators that power semantic data processing in Cortex AISQL. These operators extend SQL with natural-language semantics, allowing users to express complex semantic tasks directly within queries. For each operator, we specify its signature, behavior, and typical usage, and show how the operators compose with standard SQL while exposing data-parallel execution to support rich analytical workflows. Table 1 shows the definitions of the key semantic operators implemented in AISQL.
3.1. Map
AISQL supports simple map or projection operations using the operator. Users can use to transform text expressions using a low-level interface to an LLM and receive a text response for each row provided based on the task instruction . In addition, a PROMPT object provides an ease-of-use interface to include image or other multimodal data to the operator.
For example, users may query the system as follows:
Example result:
| AI_COMPLETE |
| This review indicates moderate dissatisfaction. |
| This review expresses positive sentiment. |
3.2. Filter
is a boolean operator, enabling usage in a SQL WHERE clause. For example, users can filter the rows of a table using a natural language predicate as follows:
3.3. Join
Using the PROMPT object, the can be extended to multiple table and column arguments for use in a SQL JOIN. For example, users can join multiple tables using natural language as follows:
A naive implementation applies AI_FILTER invocations joining tables with and rows respectively, though optimization techniques help reduce this growth rate in practice (see Section 5.3).
3.4. Classify
The operator projects each row into a discrete category selected from a finite candidate set based on a natural language instruction . In contrast to , which freely generates text, constrains the output to one of the predefined categories , effectively performing a supervised classification in natural language.
Similarly to , each classification is computed independently, supporting distributed execution. The output column may also be used as an ordinary categorical attribute in downstream SQL operations such as GROUP BY, FILTER, or AGG clauses. For example, users may compute aggregate metrics per sentiment category as follows:
This function generalizes beyond sentiment analysis to any classification problem expressible in natural language—such as intent detection, topic tagging, severity rating, or customer status identification—while maintaining compatibility with SQL’s composable semantics.
3.5. Reduce
AISQL supports two aggregate functions, AI_SUMMARIZE_AGG and that reduce a column of text values to a single aggregate result. For example, users can summarize the values of a string column as follows:
Unlike row-level operators, aggregation presents a unique challenge: tables and columns often contain more text than can fit in a single model’s context window. To address this, we employ a hierarchical aggregation strategy that processes data in batches and recursively combines intermediate results.
Our approach relies on three distinct phases:
Extract key information from a subset of the text rows into intermediate states
Recursively combine intermediate states, discarding extraneous information while synthesizing similar data
Summarize the combined state in a style and tone appropriate for the user’s context
As can be seen in the pseudo-code of Algorithm 1, the aggregate function iterates over rows, accumulating text rows in a row buffer . Once the row buffer has exceed the BATCH_SIZE token limit, generates intermediate states containing the most important information for summarization, which are all inserted into an intermediate state buffer .
Once the state buffer has exceeded BATCH_SIZE, synthesizes as many intermediate states as possible (based on the context window limit) into combined intermediate states. This process repeats until is sufficiently small.
Finally, after all rows are processed, LLM.Extract and LLM.Combine are executed until and . The final user-facing summary is generated by an LLM invocation to
3.5.1. Task Specific Aggregation
Algorithmically, is nearly identical to AI_SUMMARIZE_AGG, excepting that each LLM request to LLM.Extract, LLM.Combine, and LLM.Summarize support an additional argument for a user-provided natural language task instruction .
The user can provide a more specific task instruction for the aggregation as follows:
3.6. Supporting Multimodal Input
To support multimodal input such as images, audio, or documents, AISQL added a new data type, termed FILE. A FILE value stores a URI as well as various metadata (e.g. size, mime type, creation date, etc) of a file that lives in cloud storage and is managed by the user.
Several utility functions are provided that allow users to manage and process FILEs. For instance, processing only the image files that are stored in a Snowflake table can be performed as follows using the boolean FL_IS_IMAGE function that checks if a file is of a known image file type:
4. Customer Workloads
As discussed earlier, executing semantic operations inside a database engine introduces several system-design challenges. At the same time, customers are adapting to the AI-centric data management era and explore new workloads and usage patterns to get the most value out of AI. This section analyzes AISQL workloads executed over a three-month period (July–September 2025). We considered actual Snowflake customer workloads across multiple deployments and cloud providers. We report observations intended to inform practitioners and motivate further research. Our main observations from the workload analysis are as follows:
AI operators dominate AISQL query cost. This observation follows from the analysis of AISQL query composition by statement type (see Figure 2) and the cost distribution between LLM GPU credits and SQL Warehouse credits for each statement type (see Figure 4). Thus, in the next section, we present optimization techniques that reduce the cost of running AISQL queries with (a) AI-Aware Query Optimization that takes into account the high cost of executing AI operators to create an effective query plan (see Section 5.1); and (b) Adaptive Model Cascades that uses a lightweight proxy LLM to process most rows while using a larger oracle LLM to handle difficult cases where the proxy model is uncertain (see Section 5.2).
Semantic joins are prevalent in customer workloads.
Nearly of AISQL queries involve multiple tables (see Figure 5), and these multi-table queries account for over of total execution time (see Figure 3). Thus, we propose a new method for optimizing semantic joins that reduces the cost and latency of the operation (see Section 5.3).
5. AISQL Query Execution Engine
In this section, we describe the modifications made to Snowflake’s core database engine to enable AISQL queries, i.e. SQL queries that use one or more AI operators. Section 5.1 presents query plan optimizations tailored to AISQL workloads. Section 5.2 discusses the use of model cascade techniques to improve the efficiency of the AI_FILTER operator. Section 5.3 details an optimization that rewrites semantic joins into a multi-label classification problem using AI_CLASSIFY, and Section 5.4 covers enhancements applied to aggregation operators such as AI_AGG and AI_SUMMARIZE_AGG.
5.1. Optimizing AI Operators
One of the key challenges in optimizing AISQL queries is that AI operators are typically treated as black boxes by the optimizer. For example, without historical data from previous invocations on the same data, it is not possible to estimate the selectivity and hardware cost of an AI_FILTER predicate. At the same time, the per-row monetary and runtime costs of AI operators are typically orders of magnitude higher than those of other typical SQL operators and functions. Consequently, the suboptimal placement of an AI operator in the execution plan can have devastating effects on both cost and execution time. This problem is not new in database systems. Optimizing user-defined functions (UDFs) in database engines shares many similarities with the problem of optimizing AI operators, and many of the techniques published 20 years ago remain relevant (hellerstein6caching; hellestein93predicate; chaudhuri99optimization).
To better illustrate the challenges related to optimizing AISQL queries, let us consider the following scenario: a web application like arxiv.org manages research papers and allows users to search for papers using semantic operations. For simplicity, we assume that the content of each research paper is stored in two relational tables papers and paper_images. The former stores the basic information for each paper, such as ”title” and ”authors”, whereas the latter stores information about extracted images and graphs for each paper. Both tables have a column of type FILE that links each row with a file stored in cloud storage (e.g., AWS S3). A simplified schema and some sample data are shown in Figure 6.
Now let us consider a user who is interested in identifying the research papers that: a) were published in a specific time period, b) discuss ”energy efficiency in database systems”, and c) include a figure with results using the well-known TPC-H benchmark. Finally, the user is interested in compiling a summary of the corresponding abstracts. This query requires inspecting both the textual and image contents of research papers. Today, the user can write the following AISQL query in Snowflake as follows:
The AISQL query in this example has three predicates with different cost and performance characteristics. The first is a predicate on a ”date” column, which is relatively ”cheap” to evaluate and for which most database engines can do a reasonably good job at estimating its selectivity. The second predicate is an AI_FILTER on a VARCHAR column, which is more expensive as it invokes an LLM for each row and its selectivity is unknown at compile time. The execution cost can be estimated based on the average number of tokens in the column values. The third predicate is also an AI_FILTER, but it is applied to images and invokes a multimodal model. Multimodal models are typically larger and more expensive than models operating only on text input.
Suppose the cost of each predicate is ignored. In that case, the optimizer may generate Plan A that is shown in Figure 7, in which all predicates are ”pushed” below the join, with the underlying assumption that join is an expensive operator and hence the number of rows sent to this operator should be minimized. In addition, the most selective predicate (AI_FILTER in this example) is applied first on the rows scanned from the ”papers” table. Plan A results in 110,000 LLM calls; the numbers in the parentheses indicate the output cardinality of each operator.
Figure 7 shows a better execution plan (Plan B) that takes into account the cost of AI predicates and tries to minimize the number of LLM calls. By ”pulling” the image AI_FILTER predicate above the join and by changing the order in which predicates are applied on the rows from the ”papers” table, the compiler generates an execution plan that results in 330 LLM calls, which results in a 300 improvement in both cost and execution time!
This example highlights two of the optimizations that the Snowflake execution engine is applying to improve the performance of AISQL queries (see Section 6 for an experimental evaluation), namely: a) predicate reordering, b) optimizing query plans for AI inference cost. Predicate reordering has both a compiler and an execution component. During compilation, the predicate evaluation order is determined based on the relative cost of each predicate, i.e., the most expensive predicates are evaluated last. During execution, runtime statistics are collected on the cost and selectivity of predicates. These statistics are used to determine at runtime whether a different predicate evaluation order is more effective. For instance, if an AISQL query evaluates two AI_FILTER predicates on text columns from the same input, we can determine at runtime which one is more selective and change the evaluation order so that the more selective AI_FILTER predicate is applied first.
As we discussed in Section 4, AISQL queries typically spend a significant fraction of execution time in AI operators. Consequently, optimizing AISQL queries requires us to rethink the optimization criteria used in the compiler. We initially considered how AI_FILTER predicates are placed in the execution plan with respect to joins and aggregations, as this problem showed up in numerous customer workloads. Although the selectivity of an AI_FILTER is unknown at compile time, we identified that simply optimizing the total number of AI inference calls produces good results in many cases. However, as AISQL queries become more complex and compiler estimations less reliable, these kinds of optimizations become less effective in practice. To that end, we are considering dynamic optimization techniques and caching runtime statistics that can help the compiler make more informed decisions in subsequent AISQL queries. Additionally, we are exploring methods to generalize the compiler’s optimization of AI functions and operator placement in complex execution plans.
5.2. Adaptive Model Cascades
AI operators such as AI_FILTER must often process millions of rows in production database workloads. However, invoking an LLM on every row introduces prohibitive costs and latency. To address this challenge, the AISQL engine implements model cascades: a lightweight proxy model processes most rows, while a powerful oracle model handles only difficult cases where the proxy exhibits uncertainty. The proxy model’s predictions are accepted when confidence is high, but uncertain cases are escalated to the oracle for more reliable decisions.
Implementing model cascades in a production database system presents unique challenges. First, the system must learn an effective routing strategy online from streaming data without prior knowledge of the query’s data distribution. Second, the algorithm must operate in a highly concurrent environment, where multiple workers process different portions of the data in parallel. Third, it must honor user-specified cost budgets while maintaining high-quality results. Finally, the routing decisions must adapt dynamically as data characteristics change across batches.
Snowflake’s AISQL engine addresses these challenges through an adaptive threshold learning algorithm that reframes cascade routing as a statistical estimation problem. The algorithm uses a sample of oracle evaluations to learn confidence thresholds that partition rows into three regions: an accept region where the proxy is confident and predicts positive, a reject region where proxy confidence is low and predicts negative, and an uncertainty region that might require invoking the oracle and making conservative decisions.
The algorithm operates in four phases over batches of rows:
Phase 1: Proxy Inference. Execute the proxy model on all rows in the current batch to obtain predictions and confidence scores. For an AI_FILTER predicate, each row receives a score representing the proxy’s confidence that the predicate is satisfied.
Phase 2: Sampling. Allocate a fraction of the oracle budget to sample rows for ground-truth labels. Rather than sampling uniformly, the algorithm uses importance sampling to oversample rows with intermediate confidence scores. These are the rows where the proxy is most uncertain and where oracle feedback provides maximum information for threshold estimation.
Phase 3: Threshold Learning. Using the sampled oracle labels, compute two thresholds and that define the uncertainty region:
Rows with are predicted negative (proxy is confident it’s false)
Rows with are predicted positive (proxy is confident it’s true)
Rows with remain in the uncertainty region
These thresholds are computed using concentration inequalities that provide probabilistic guarantees with limited samples (kang2020approximate). The algorithm accounts for sampling bias through importance weighting, ensuring threshold estimates are unbiased despite non-uniform sampling.
Phase 4: Adaptive Refinement. The algorithm refines thresholds iteratively across batches. For each new batch: (1) identify rows falling in the current uncertainty region, (2) allocate remaining oracle budget to sample additional rows from this region, (3) update threshold estimates using the new sample, and (4) repeat until the oracle budget is exhausted or the uncertainty region is empty. Iterative refinement allows the thresholds to adapt to evolving data distributions.
Once thresholds are learned, the engine makes final routing decisions. Rows outside the uncertainty region are classified directly using learned thresholds and proxy predictions. Rows remaining in the uncertainty region are routed to the oracle if budget permits; otherwise, the proxy prediction is used as a fallback.
The iterative batch-based implementation in the AISQL engine naturally accommodates concurrent database execution. Multiple workers process different data partitions in parallel, each collecting oracle samples within their assigned batches. These samples may be periodically aggregated to update shared threshold estimates, allowing the algorithm to learn from the entire data stream while workers proceed independently.
5.3. Query Rewriting for Semantic Joins
Semantic joins using AI_FILTER enable powerful natural language-based table joining capabilities. For example, let us consider a scenario where a user has two Snowflake tables: (1) “Reviews” contains product reviews, and (2) “Categories” includes names of product categories. A simplified schema of these two tables, along with some sample data, is shown in Figure 8.
Consider an example where a user wants to associate every product review with one or more categories. The user can write the following AISQL queries that join Reviews with Categories using an AI_FILTER predicate in the ON clause.
Due to the semantic nature of the join condition, efficient join algorithms such as hash-join or sort-merge join, which rely on equality predicates, are not applicable. The AISQL execution engine must therefore treat the semantic join as a cross join, thereby generating all possible pairs from Reviews and Categories and evaluating AI_FILTER for each pair. Therefore, many queries like this can result in thousands of LLM calls and high inference costs, even for small datasets. More specifically, the above approach requires LLM calls, where and denote the cardinalities of the respective tables.
After analyzing a wide range of customer queries, we discovered that, in many cases, semantic joins are equivalent to a multi-label classification problem. For instance, in the previous example, rather than applying AI_FILTER for every review-category pair, we can leverage the Snowflake AI_CLASSIFY function, which can classify a text or image input into one or more user-specified labels. By re-framing the semantic join into a classification problem we can instead execute AI_CLASSIFY for every row of table Reviews using the values of column Categories.label as labels, thereby reducing the number of AI inference calls from 24 (4 reviews 6 labels) to just 4 (one for each review), a 6 improvement in the example above.
Automatically detecting such cases is not a trivial task. During the compilation of AISQL queries, we introduced an AI-based oracle that examines every semantic join to determine if it can be rewritten as a multi-label classification problem. The oracle analyzes the user-specified natural language prompt, schema metadata (e.g., column and table names), statistics (e.g., number of distinct values), as well as sample values from each input source. Using all that information, the oracle determines during compilation if a semantic join should be converted into a classification task and which input dataset contains the labels for the classification. Subsequently, the transformation is performed as a regular rewrite operation by the compiler. The rewritten query is equivalent to the original user query but adds additional operations to: a) reduce the number of input labels to each AI_CLASSIFY call, and b) handle cases where the input of each AI_CLASSIFY call fits in the context window of the user-specified LLM model.
5.4. AI Aggregation
Earlier, we described an incremental fold strategy for combining intermediate states within a Reduce operation for both and AI_SUMMARIZE_AGG. However, invoking three separate LLM APIs for LLM.Extract, LLM.Combine, and LLM.Summarize introduces significant overhead for small datasets that fit within the model context window. Therefore, a simple ”short-circuit” alternative provides opportunities for throughput improvements by identifying these scenarios and skipping unnecessary incremental fold steps. This simple algorithm change measured an 86.1% latency reduction in AI_SUMMARIZE_AGG on queries using small datasets.
6. Experimental Evaluation
In this section, we present the experimental evaluation of selected techniques introduced in the preceding sections. Due to space constraints, we focus on a representative subset of results that exemplify the key findings and highlight the practical effectiveness of the proposed methods. All experiments were performed using a production-release version of Snowflake.
6.1. AI-aware Query Optimization
In this section, we evaluate the performance impact of two compiler optimizations: a) predicate reordering and b) LLM cost-optimized predicate placement with respect to joins. Figure 9 shows the performance impact of predicate reordering for an AISQL query that is shown below. The SQL statement has two predicates in the WHERE clause, an IN predicate and an AI_FILTER predicate on a VARCHAR column. We used a dataset with 1000 articles from the New York Times111https://www.nytimes.com/.
In this experiment, we vary the selectivity of the IN predicate between and , with indicating that all input rows satisfy the IN predicate. We then measure the speedup we get by carefully reordering the predicate evaluation so that the AI_FILTER predicate is evaluated last. As can be seen from the normalized results in Figure 9, applying the AI_FILTER last yields up to a 7 speedup in our experiment. In practice, the actual speedup will depend on several factors, such as the relative costs of the different predicates and their selectivities.
In the second experiment, we measure the impact on total execution time of the compiler optimization that decides the placement of AI predicates (AI_FILTER) with respect to joins based on the total LLM cost. We consider the following AISQL statement that performs a join between two input tables and has two predicates in the ON clause, one of which is an AI_FILTER.
For this experiment, we adjust the ratio of input rows from the left input (NY_ARTICLES_V1), where the AI_FILTER is applied, to the total number of output rows generated by the join. We varied the ratio between and , where means that the join produces twice as many rows as the cardinality of the left input. We compare the optimization that considers the total LLM cost when deciding the placement of AI_FILTER against two other approaches: a) Always Pull-up and, b) Always Push-down. Always Pull-up always ”pulls” AI predicates on top of joins, whereas Always Push-down is the default behavior in Snowflake’s optimizer that always ”pushes” predicates below joins.
As can be seen from the results in Figure 10, neither Always Pull-up nor Always Push-down is optimal across the different output/input ratios. When the ratio is (i.e., a selective join), Always Pull-up performs better. In contrast, Always Push-down is better when the output/input ratio is (i.e., explosive joins). Considering the total cost of LLM when deciding the placement of AI_FILTER predicates (termed AI-aware Optimization) achieves the best performance across the entire range of measured output/input ratios. In practice, the effectiveness of this technique relies on the compiler’s ability to accurately estimate join selectivity, which is not a trivial problem, especially in complicated queries with a large number of joins. Future efforts will explore adaptive optimization techniques, which involve evaluating various execution plans during compilation while collecting runtime statistics, such as the cost and selectivity of AI operators, from sample data to improve decisions regarding AI operator placement.
6.2. Adaptive Model Cascades
We evaluate the performance impact of model cascades in AI_FILTER operations across seven Boolean classification datasets. As detailed in Section 5.2, the cascading approach uses adaptive threshold learning to route predictions between a lightweight proxy model and a powerful oracle model. We compare three configurations: (1) the baseline approach using only the oracle model Llama3.3-70B for all predictions, (2) the cascade approach combining both models with adaptive threshold learning, and (3) using only the proxy model Llama3.1-8B for all predictions.
Our evaluation methodology executes each query five times to ensure statistical reliability. We measure both execution efficiency (query time) and prediction quality using standard classification metrics: accuracy, precision, recall, and F1 score. All quality assessments use ground-truth labels from the evaluation datasets as a reference.
Table 2 reports the mean performance across all datasets, including execution time, speedup relative to the baseline, F1 score with delta () relative to the baseline, and precision/recall metrics. All speedup percentages and delta F1 values are computed with respect to the baseline approach.
| Method | Time [s] | Speedup | F1 | Prec. / Rec. | |
| Score | |||||
| llama3.3-70B | 975.9 | — | 0.812 | — | 0.813 / 0.829 |
| llama3.1-8B | 296.2 | 3.3x | 0.659 | -18.8% | 0.704 / 0.686 |
| Cascade | 336.4 | 2.9x | 0.777 | -4.3% | 0.784 / 0.794 |
Performance gains. The cascade approach achieves substantial speedup with mean execution time reduced by 65.5% (from 975.9 to 336.4 seconds), corresponding to a 2.9 speedup. This improvement stems from the cascade’s ability to route the majority of predictions through the lightweight proxy model, reserving the more expensive oracle model only for uncertain cases. For comparison, the proxy model alone runs in 296.2 seconds (3.3 speedup), as expected from its smaller model size, though this comes at a substantial quality cost.
Figure 11 shows per-dataset results, revealing that speedup varies considerably across workloads. The NQ dataset achieves the highest speedup at 5.85, while QUORA and FARL show moderate improvements (1.22 and 1.30 respectively). These variations reflect differences in how the threshold learning algorithm partitions the data. Easier workloads, where the proxy model exhibits stronger confidence scores, yield higher routing rates to the proxy and thus greater speedups.
Quality analysis. The cascade maintains strong prediction quality with only modest mean degradation. The F1 score decreases by 4.3% (from 0.812 to 0.777), with balanced precision and recall remaining at 0.784 and 0.794, respectively. This demonstrates that the threshold learning logic effectively identifies rows where the proxy model’s confidence scores are reliable. In contrast, using the proxy model exclusively results in substantially degraded quality, with F1 dropping by 18.8% to 0.659, demonstrating the critical role of the oracle model in handling predictions that fall within the uncertainty region.
As shown in Figure 11, the impact on quality scores varies across datasets. NQ maintains near-identical F1 scores (), indicating that most predictions have clear confidence patterns suitable for routing. IMDB and SST2 show moderate degradation ( and respectively), while BOOLQ and QUORA experience larger drops ( and ). These variations suggest that cascade effectiveness depends on dataset characteristics, i.e., tasks where the proxy model’s uncertainty correlates well with prediction difficulty benefit most from the adaptive routing strategy.
Discussion. The cascade offers an attractive balance between the baseline and proxy-only approaches. For applications where near-optimal quality is essential, the baseline achieves the highest F1 score (0.812) but at a 2.9 higher computational cost. For latency-critical workloads where moderate quality is acceptable, the cascade delivers 65.5% speedup while retaining 95.7% of baseline F1 performance. The proxy-only configuration is suitable only when speed is paramount and substantial quality degradation is tolerable.
Beyond the speed-quality trade-off, the cascade system provides two additional control mechanisms. Users can restrict the oracle budget by specifying a maximum number of oracle invocations, ensuring predictable costs regardless of data characteristics. The system also supports precision and recall targets, allowing users to specify desired quality thresholds. However, achieving these targets may require higher delegation rates to the oracle model, which reduces the speedup gains.
We also found the cascade algorithm to be resilient to both the degree of parallelism in the SQL execution engine and the number of rows processed by individual workers.
6.3. Query Rewriting for Semantic Joins
We evaluate AI_CLASSIFY rewrite optimization using eight semantic join benchmarks that cover entity matching (ABTBUY, NASDAQ), document categorization (ARXIV, EURLEX, NYT, CNN), news classification (AG NEWS), and biomedical concept linking (BIODEX). Each benchmark consists of two tables with natural language content, where the join task is to identify semantically related row pairs. For example, in ABTBUY the predicate matches product descriptions from different e-commerce sites that refer to the same item, while in NYT it links news articles to their topical categories. The left table (denoted as in Table 4) and the right table () span 50 to 500 rows, yielding up to 250,000 candidate pairs that must be considered by the naive join approach using filtering.
| Method | Time [s] | Speedup | F1 | Prec. / Rec. | |
| Score | |||||
| Cross Join | 2330.56 | — | 0.412 | — | 0.388 / 0.761 |
| Classify Rewrite | 40.96 | 30.7x | 0.596 | +44.7% | 0.745 / 0.540 |
Table 3 summarizes the mean performance across all benchmarks. Table 4 reports the number of LLM calls, the total execution time (seconds), and standard classification metrics (precision, recall, F1) for both approaches across datasets of varying cardinalities.
Table 4 reports the number of LLM calls, the total execution time (seconds), and standard classification metrics (precision, recall, F1) for both approaches across datasets of varying cardinalities. Performance gains. The query rewriting achieves dramatic speedups across all benchmarks, ranging from (ARXIV) to (CNN), with a mean speedup of . These improvements stem directly from reducing LLM invocations: the baseline requires calls, while the rewrite needs only classifications. For the CNN dataset with 500 rows per table, this reduces 250,000 binary classifications to just 1,000 multi-label classifications, cutting execution time from 4.4 hours to 3.8 minutes. The speedup magnitude correlates strongly with dataset scale; larger Cartesian products yield proportionally greater benefits.
| Dataset | L | R | Cross Join (AI_FILTER) | AI_CLASSIFY Rewrite | Speed | ||||||||
| Calls | Time | Precision | Recall | F1 | Calls | Time | Precision | Recall | F1 | ||||
| NASDAQ | 100 | 100 | 10000 | 51.46 | 0.029 | 0.96 | 0.056 | 100 | 2.69 | 0.851 | 0.731 | 0.788 | |
| EURLEX | 50 | 194 | 9700 | 39.93 | 0.172 | 0.833 | 0.286 | 50 | 2.37 | 0.86 | 0.21 | 0.338 | |
| BIODEX | 50 | 197 | 9850 | 79.6 | 0.118 | 0.585 | 0.195 | 50 | 2.81 | 0.409 | 0.2 | 0.269 | |
| ABTBUY | 100 | 100 | 10000 | 60.27 | 0.967 | 0.967 | 0.967 | 100 | 2.51 | 0.968 | 0.968 | 0.968 | |
| AG NEWS | 100 | 100 | 10000 | 63.21 | 0.565 | 0.87 | 0.685 | 100 | 2.57 | 0.91 | 0.61 | 0.731 | |
| AG NEWS | 200 | 200 | 40000 | 192.27 | 0.505 | 0.8 | 0.619 | 200 | 6.08 | 0.905 | 0.61 | 0.728 | |
| ARXIV | 500 | 500 | 250000 | 646.67 | 0.55 | 0.18 | 0.27 | 1500 | 42.31 | 0.549 | 0.2 | 0.293 | |
| NYT | 500 | 500 | 250000 | 1618.58 | 0.034 | 0.775 | 0.065 | 1500 | 39.39 | 0.609 | 0.414 | 0.493 | |
| CNN | 500 | 500 | 250000 | 15955.67 | 0.729 | 0.99 | 0.84 | 1000 | 229.48 | 0.807 | 0.984 | 0.887 | |
Quality analysis. The impact of rewrite optimization on prediction quality varies by dataset characteristics. On entity matching tasks with clear semantic signals (ABTBUY), both approaches achieve nearly identical F1 scores ( vs ), demonstrating that the rewrite preserves quality when matches are unambiguous. More remarkably, the rewrite substantially improves quality on datasets where the baseline suffers from poor precision. For NASDAQ, the extremely low precision of the baseline () results in an F1 score of , while the rewrite achieves a precision of and an F1 score of . We attribute this to the multi-label classification paradigm: by presenting all candidate labels simultaneously, AI_CLASSIFY enables better comparative reasoning than isolated binary decisions in AI_FILTER. The mean F1 improves by (from to ), with precision increasing from to . Overall, the mean F1 improves by (from to ), with precision increasing from to .
Trade-offs. The rewrite exhibits precision-recall trade-offs that depend on the classification strategy. Datasets like EURLEX and BIODEX show recall degradation ( and , respectively) despite improved precision. Recall degradation occurs when the model conservatively selects matches from the full set of labels, prioritizing accuracy over coverage. For applications where recall is critical, hybrid strategies that combine both approaches or tune the classification prompt may be necessary.
7. Related Work
The intersection of databases and AI has evolved along two directions: AI4DB, which utilizes ML to enhance database systems themselves, and DB4AI, which integrates native AI capabilities into databases to simplify analytics and model inference. AISQL builds upon both paradigms by integrating LLMs directly into the SQL execution layer.
The AI4DB line of research applies ML algorithms to internal database components such as query optimization, indexing, and configuration tuning. Early systems like LEO (stillger2001leo) used feedback learning to refine cost estimates for the DB2 optimizer. Later, OtterTune (vanaken2017ottertune) and CDBTune (zhang2019end) leveraged large-scale supervised and reinforcement learning to automatically tune database parameters, inspiring the vision of self-driving databases (pavlo2017self).
A rich body of work explores learned query processing components. Learned Indexes (kraska2018case) replaced traditional index structures with neural networks, while ALEX (ding2020alex) extended these ideas to support dynamic workloads. In query optimization, learned cardinality estimation models, such as MSCN (kipf2018learned), Naru (yang2019deep), and DeepDB (hilprecht2020deepdb), achieved notable improvements in selectivity prediction by learning correlations across attributes and tables. Reinforcement learning has also been applied to plan search and join ordering, exemplified by ReJOIN (marcus2018deep), Neo (marcus2019neo), and Bao (marcus2021bao). These works culminate in the SageDB (kraska2019sagedb), which proposes a learned database system where each component is automatically specialized for a given data distribution and workload.
The application of ML to query optimization extends to handling expensive predicates in specialized domains. Systems such as BlazeIt (kang2019blazeit) and TASTI (kang2020tasti) optimize queries involving expensive ML-based predicates using model cascades, semantic indexes, and query-specific proxy models. In particular, model cascades that combine cheap approximate models with expensive accurate ones (kang2020approximate; chen2023frugal; zellinger2025rational) have influenced the design of AISQL’s approach to managing expensive AI operations.
The DB4AI perspective takes the inverse view: enabling databases to host and execute ML models or reasoning workloads natively, rather than treating them as external applications. This line of work seeks to unify analytics and model inference under the relational abstraction, with the goal of allowing users to express AI tasks declaratively using SQL.
Earlier in-database ML frameworks such as MADLib (hellerstein2012madlib) extended SQL with support for descriptive statistics and simple ML methods. Recent systems go further by integrating semantic and language-model capabilities directly into query processing, for example LOTUS (patel2025semantic) introduces a set of semantic operators for Pandas-like dataframe processing. ThalamusDB (jo2025thalamusdb) explores approximate query processing for multimodal data, supporting natural language predicates over visual, audio, and textual content through zero-shot models combined with relational operators.
Several systems have proposed specialized approaches to LLM-powered SQL processing. SUQL (liu2024suql) augments SQL with answer and summary operators for knowledge-grounded conversational agents, with the focus on row-wise LLM operations for question answering. ZenDB (lin2024towards) optimizes SQL queries for extracting structured data from semi-structured documents with predictable templates. Palimpzest (liu2025palimpzest) provides a declarative framework for LLM-powered data processing with specialized map-like operations and basic cost-based optimizations. UQE (dai2024uqe) studies embedding-based approximations for LLM-powered filters and stratified sampling for aggregations, though it provides best-effort performance without accuracy guarantees.
Beyond SQL-centric approaches, several frameworks have emerged for LLM-powered data processing. Aryn (anderson2024design) offers a Spark-like API with PDF extraction and human-in-the-loop processing capabilities. DocETL (shankar2025docetl) proposes agent-driven pipeline optimization for complex document processing tasks and uses LLM agents to explore task decomposition strategies. EVAPORATE (arora2023language) specializes in extracting semi-structured tables through code synthesis. These systems demonstrate the potential of LLM-powered processing, but often lack formal cost models and provide limited performance guarantees.
Optimizing expensive predicates and user defined functions was the subject of a number of research papers in the past (hellerstein6caching) (hellestein93predicate) (chaudhuri99optimization). This line of work inspired many of the optimizations performed in AISQL. In (hellestein93predicate) a cost-based optimization technique was introduced that improved query performance by selectively reordering or “migrating” expensive selection predicates in relational query plans. The authors developed an analytical model and algorithm that balances predicate selectivity and evaluation cost, achieving faster query execution without exhaustive reordering. (hellerstein6caching) presents strategies to optimize query performance by caching the results of costly user-defined or external function calls within a query. It introduces algorithms that determine when and how to reuse previously computed method results, reducing redundant evaluations and significantly improving overall query efficiency. Finally, (chaudhuri99optimization) proposed techniques to integrate expensive, user-defined predicates into traditional cost-based query optimization frameworks. It developed a formal model for estimating predicate costs and selectivities, enabling the optimizer to reorder and combine such predicates effectively to minimize overall query execution time.
AISQL unifies the goals of both research directions. From AI4DB, it borrows the principle of learning to optimize workloads by applying cost modeling, feedback, and adaptive execution to manage expensive semantic operators. From DB4AI, it inherits the vision of integrating LLM-powered inference directly into the SQL engine. Unlike systems that provide best-effort performance or focus solely on row-wise operations, AISQL co-optimizes query execution and LLM inference within a unified relational framework. AISQL represents a step toward production-ready AI-native database systems that reason, learn, and optimize within the same execution environment.
8. Conclusions
The proliferation of unstructured data alongside traditional structured datasets has created a fundamental gap in modern data warehouses: users need to perform semantic reasoning over hybrid data but are forced to export data to external systems, orchestrate complex pipelines, and sacrifice the optimization benefits of integrated query processing. This paper presents Snowflake AISQL, a production SQL engine that bridges this gap by integrating LLM capabilities directly into SQL through semantic operators.
AISQL introduces six semantic operators that compose naturally with relational SQL primitives. These operators enable users to express complex analytical tasks that blend structured queries with semantic reasoning in a single declarative statement, and eliminate the need for external data movement and manual pipeline orchestration.
The core technical challenge we address is making semantic operators efficient at a production scale. LLM-based semantic operators are orders of magnitude more expensive than traditional SQL operations; their cost and selectivity are opaque to the query optimizer, and conventional optimization heuristics can produce plans that are catastrophically inefficient. Our solution comprises three techniques driven by analysis of real customer workloads. AI-aware query optimization treats LLM inference cost as a first-class objective, resulting in efficient operator placement and predicate evaluation order. Adaptive model cascades reduce costs through intelligent routing between lightweight proxy models and powerful oracle models. Query rewriting for semantic joins transforms quadratic-complexity joins into a linear multi-label classification task and, as a result, delivers speedups with improved prediction quality.
Our production deployment at Snowflake validates these design choices. Analysis of customer workloads reveals that AI operators consume the majority of execution costs, and nearly half of queries involve multi-table operations. These observations directly motivate our focus on join optimization and cost-aware planning. Extensive experimental evaluation across diverse benchmarks demonstrates that AISQL successfully delivers both efficiency and quality at the production scale.
AISQL represents a step toward unified data processing architectures where semantic reasoning over unstructured content is a first-class database capability rather than an external add-on. By integrating LLM capabilities into the SQL engine, with a focus on cost optimization, we enable organizations to unlock insights from their data assets through familiar declarative interfaces, while maintaining the performance, scalability, and reliability expected from modern data warehouses.